# INR-TH-2025-001

### v2

@@ Abstract
with $h_{eff}$ referring to the effective number of degrees of freedom in the cosmic plasma entering the entropy density of the interesting epoch. The sterile neutrino production for the chosen model parameters terminates at about $T=30$ MeV, and from that moment we run the FortEPiaNO code with initial value of $h_{eff}(30\,\text{MeV})=12.47$ [22] and initial neutrino asymmetries $\Delta_{\nu_{\alpha}}$ (the outcome of the dmpheno running) equal to $(5.97,-3.01,-2.95)\cdot 10^{-4}$ (left panel) and $(87.9707,-46.9718,-35.6134)\cdot 10^{-4}$ (right panel).
@@
4. There is more room for mixing with muon or tau neutrinos, where the initial electron neutrino asymmetry is close to zero, and muon and tau neutrino asymmetries are large, equal in value but of opposite sign. Such initial conditions may be obtained with specific symmetry in the sector responsible for generation of the initial lepton asymmetry in the primordial plasma (possibly after electroweak transition). As we discussed above, the active neutrino oscillations between muon and tau neutrinos start earlier, at 3 times higher temperature, and the cancellation of the total asymmetry in this case is much more efficient. In Fig. 4
@@
Figure 4: Sterile neutrino spectrum at various temperatures (left panel) and subsequent evolution of the neutrino flavor asymmetries (right panel), electron neutrino (red lines), muon neutrino (green lines) and tau neutrino (blue lines) with temperature at the oscillation stage. The initial lepton asymmetries at 20 MeV are $(-6.0208,128.8341,-106.8030)\cdot 10^{-4}$, sterile neutrino mass $m=7.11$ keV, the mixing with muon neutrino is tuned to fully explain the dark matter component, $\sin^{2}2\theta=1.84\cdot 10^{-14}$. The calculations are performed for the best fit parameters of the neutrino sector taken from Ref. [16] for both normal (solid lines) and inverse (dotted lines) hierarchies.
@@
we illustrate this point by numerical calculations performed for the mixing with muon neutrino, $\sin^{2}2\theta=3\cdot 10^{-14}$, $m_{\nu_{s}}=7.11$ keV and the initial set of neutrino asymmetries $(0,45,-45)\cdot 10^{-4}$, which evolves to $(-6.0208,128.8341,-106.8030)\cdot 10^{-4}$ by the onset of active neutrino oscillations. As we discussed above, see Fig. 1, with higher lepton asymmetry the sterile neutrino production terminates later, so we run the FortEPiaNO code from $T=20$ MeV when $h_{eff}=11.28$. One observes that the muon and tau lepton asymmetries equalize very rapidly and neatly, as expected, reaching values $309.90\cdot 10^{-4}$ and $309.93\cdot 10^{-4}$ at $T=6.9$ MeV. We find also that the asymmetry dynamics is somewhat different in the cases of normal and inverse mass hierarchy, see right panel of Fig. 4.
@@
Note, that while we start with zero total neutrino asymmetry, it becomes non-zero after the sterile neutrino generation, since the asymmetry, mostly from the flavor which mixes with sterile neutrinos, gets transferred to the sterile neutrinos. Some asymmetry also emerges in active neutrinos, which do not mix with the sterile component. In principle, one can tune the initial pattern of neutrino flavor asymmetry in such a way, that after termination of the sterile neutrino production the electron neutrino asymmetry equals zero. It requires some small but non-zero initial asymmetry in the electron neutrinos. Likewise, one can start with a set of sterile neutrino asymmetries, when after production both the electron neutrino and the total lepton asymmetries are zero. Whether such tunings may be adopted within concrete mechanisms capable of generating the neutrino asymmetries in the early Universe remains to be explored.
@@
5. To summarize, in this letter we show that the presently allowed model parameter space of the sterile neutrino dark matter may be significantly extended with specific tuned neutrino flavor asymmetries. If large in magnitude, but the sum is close to zero, it can be efficiently cancel in the subsequent oscillations of the active neutrino. The inverse ordering of neutrino masses shows more efficient cancellation. Hence, at the moment of neutron freeze-out the electron asymmetry may be greatly reduced, even to zero. It mitigates the limits from primordial nucleosynthesis, allows for higher initial lepton asymmetry and hence lower sterile-neutrino mixing angles. Likewise, with a higher initial lepton asymmetry the resonant mechanism produces colder sterile neutrinos. We show, that if the lepton asymmetry disappears from the plasma, the resulting spectrum is indeed colder. At the same time, higher asymmetry postpones the sterile neutrino production, and with sufficiently high asymmetry the active neutrino oscillations may do the job, cancel the asymmetry and make the sterile neutrino spectrum cold.
@@
Without using a particular mechanism to produce the original lepton asymmetry, one is quite free to tune the parameters in the sterile neutrino sector. We illustrate our statements with numerical calculations performed for particular values of sterile neutrino mass and mixing angle. The parameters of the active neutrino sector are fixed by neutrino oscillation experiments, except the mass hierarchy and the value of CP-phase. As we observed above, the inverted hierarchy is capable of reducing the net neutrino asymmetry more efficiently. We fix the value of CP-phase as suggested in [16], but the present accuracy of it’s determination is poor. Thus we run the FortEPiaNO code with the suggested value of CP-phase and with zero CP-phase for the set of neutrino asymmetries and do not find any significant change in the dynamics of asymmetry, see Fig.,5.
@@
Figure 5: Evolution of the neutrino flavor asymmetries for the normal hierarchy (left panel), and inverse hierarchy (right panel), with taken from [16] values of CP-phase (dotted lines) and zero CP-phase (solid lines). The initial pattern of the lepton asymmetries for both plots are the same as for the left plot in Fig. 3.

### v3

@@ Abstract
with $h_{eff}$ referring to the effective number of degrees of freedom in the cosmic plasma entering the entropy density of the interesting epoch. The sterile neutrino production for the chosen model parameters terminates at about $T=30$ MeV, and from that moment we run the FortEPiaNO code with initial value of $h_{eff}(30\,\text{MeV})=12.47$ [23] and initial neutrino asymmetries $\Delta_{\nu_{\alpha}}$ (the outcome of the dmpheno running) equal to $(5.97,-3.01,-2.95)\cdot 10^{-4}$ (left panel) and $(146.61,-78.29,-59.36)\cdot 10^{-4}$ (right panel). We use linearly spaced 30-point momentum grid from $y_{min}=0.01$ to $y_{max}=20$ with full momentum dependency. Such settings ensure a sufficient accuracy of the calculations. To perform this calculation, we modified the public version of the code FortEPiaNO by introducing the neutrino asymmetry matrix, which accounts for the neutrino asymmetry potential $V_{\alpha}^{L}$.
@@
4. So far we considered the sterile neutrino mixed with electron neutrino. There is more room for mixing with muon or tau neutrinos, where the initial electron neutrino asymmetry is close to zero, and muon and tau neutrino asymmetries are large, equal in value but of opposite sign. Such initial conditions may be obtained with specific symmetry in the sector responsible for generation of the initial lepton asymmetry in the primordial plasma (possibly after electroweak transition). As we discussed above, the active neutrino oscillations between muon and tau neutrinos start earlier, at 3 times higher temperature, and the cancellation of the total asymmetry in this case is much more efficient. In Fig. 4
@@
Figure 4: Sterile neutrino spectrum at various temperatures (left panel) and subsequent evolution of the neutrino flavor asymmetries (right panel), electron neutrino (red lines), muon neutrino (green lines) and tau neutrino (blue lines) with plasma temperature, $x\equiv m_{e}/T$, at the oscillation stage. The initial lepton asymmetries at 20 MeV are $(-6.0208,128.8341,-106.8030)\cdot 10^{-4}$, sterile neutrino mass $m=7.11$ keV, the mixing with muon neutrino is tuned to fully explain the dark matter component, $\sin^{2}2\theta=7.36\cdot 10^{-14}$. The calculations are performed for the best fit parameters of the neutrino sector taken from Ref. [16] for both normal (solid lines) and inverse (dotted lines) hierarchies.
@@
we illustrate this point by numerical calculations performed for the mixing with muon neutrino, $\sin^{2}2\theta=1.2\cdot 10^{-13}$, $m_{\nu_{s}}=7.11$ keV and the initial set of neutrino asymmetries $(0,45,-45)\cdot 10^{-4}$, which evolves to $(-6.0208,128.8341,-106.8030)\cdot 10^{-4}$ by the onset of active neutrino oscillations. As we discussed above, see Fig. 1, with higher lepton asymmetry the sterile neutrino production terminates later, so we run the FortEPiaNO code from $T=20$ MeV when $h_{eff}=11.28$. One observes that the muon and tau lepton asymmetries equalize very rapidly and neatly, as expected, reaching values $309.90\cdot 10^{-4}$ and $309.93\cdot 10^{-4}$ at $T=6.9$ MeV. We find also that the asymmetry dynamics is somewhat different in the cases of normal and inverse mass hierarchy, see right panel of Fig. 4.
@@
Note, that while we start with zero total neutrino asymmetry, it becomes non-zero after the sterile neutrino generation, since the asymmetry, mostly from the flavor which mixes with sterile neutrinos, gets transferred to the sterile neutrinos. Some asymmetry also emerges in active neutrinos, which do not mix with the sterile component. In principle, one can tune the initial pattern of neutrino flavor asymmetry in such a way, that after termination of the sterile neutrino production the electron neutrino asymmetry equals zero. It requires some small but non-zero initial asymmetry in the electron neutrinos. Likewise, one can start with a set of sterile neutrino asymmetries, when after production both the electron neutrino and the total lepton asymmetries are zero. Whether such tunings may be adopted within concrete mechanisms capable of generating the neutrino asymmetries in the early Universe remains to be explored. It is worth mentioning that with a mass in the keV range the sterile neutrino itself contributes to the energy density at BBN. Its number density at present is of about 1/cm3 or lower. Comparing it to about 100/cm3, the present number density of the active neutrino specie, see e.g. [26], one concludes that the sterile neutrino contribution to the energy density in the BBN epoch is less than one per cent of that of a single active neutrino specie, and hence is negligible, given the present and near future accuracy expected in measurements of the primordial abundances of the light chemical elements, see e.g. [16].
@@
5. To summarize, in this letter we show that the presently allowed model parameter space of the sterile neutrino dark matter may be significantly extended with specific tuned neutrino flavor asymmetries. If large in magnitude, but the sum is close to zero, it can be efficiently canceled in the subsequent oscillations of the active neutrino. The inverse ordering of neutrino masses shows more efficient cancellation. Hence, at the moment of neutron freeze-out the electron asymmetry may be greatly reduced, even to zero. It mitigates the limits from primordial nucleosynthesis, allows for higher initial lepton asymmetry and hence lower sterile-neutrino mixing angles. Likewise, with a higher initial lepton asymmetry the resonant mechanism produces colder sterile neutrinos. We show, that if the lepton asymmetry disappears from the plasma, the resulting spectrum is indeed colder. At the same time, higher asymmetry postpones the sterile neutrino production, and with sufficiently high asymmetry the active neutrino oscillations may do the job, cancel the asymmetry and make the sterile neutrino spectrum cold.
@@
Without using a particular mechanism to produce the original lepton asymmetry, one is quite free to tune the parameters in the sterile neutrino sector. We illustrate our statements with numerical calculations performed for particular values of sterile neutrino mass and mixing angle. The parameters of the active neutrino sector are fixed by neutrino oscillation experiments, except the mass hierarchy and the value of CP-phase. As we observed above, the inverted hierarchy is capable of reducing the net neutrino asymmetry more efficiently. We fix the value of CP-phase as suggested in [16], but the present accuracy of it’s determination is poor. Thus we run the FortEPiaNO code with the suggested value of CP-phase and with zero CP-phase for the set of neutrino asymmetries and do not find any significant change in the dynamics of asymmetry, see Fig. 5,
@@
Figure 5: Evolution of the neutrino flavor asymmetries with plasma temperature, $x\equiv m_{e}/T$, for the normal hierarchy (left panel), and inverse hierarchy (right panel), with taken from [16] values of CP-phase (dotted lines) and zero CP-phase (solid lines). The initial pattern of the lepton asymmetries for both plots is the same as for the left plot in Fig. 3.
that is consistent with detailed studies [28, 29] of the impact of the CP-phase on the evolution of neutrino flavor asymmetries.